OpenStudy (anonymous):

Help please. A car rental company increases daily rental fees 15% in the summer to cover increased fuel costs. They then have a “25% off” promotion for the fall. If a car rented for $36.00 per day before the summer, what would the per-day rental cost be during the fall promotion? Show all of your steps, each on a new line, with an explanation of why you performed that operation next to the step.

OpenStudy (anonymous):

"'15% of 36 = 15/100*36 = 5.40 new rental fees = 36+5.40=41.40 25% discount = 25/100*41.40 = 10.35 per-day rental cost be during the fall promotion = 41.40-10.35 = 31.05 let principal = x interest = x*7.5/100*5=0.375x amount=x+0.375x=1.375'x=76.50 x=76.50/1.375=55.6363"